OpenStudy (anonymous):

the ratio of females to males on a particular flight was 2 : 3. Females represented five more than 1/3 of all people aboard. how many people were on that flight

myininaya (myininaya):

i think i just gave you the two equations and you deleted it

OpenStudy (anonymous):

im so sorry i guess i didnt see it

OpenStudy (anonymous):

A15 b.25 c.30 d40 e75

myininaya (myininaya):

\[\frac{f}{m}=\frac{2}{3}\] \[f=5+\frac{1}{3}(f+m)\] do you see how i got these two equations?

OpenStudy (anonymous):

yeah i see

myininaya (myininaya):

ok solve the first equation for f by multiplying both sides by m then substitute that f into the second equation and solve for m

OpenStudy (anonymous):

how?

myininaya (myininaya):

have you solve the first equation for f yet?

OpenStudy (anonymous):

im not sure how to get f

myininaya (myininaya):

\[\frac{f}{m}=\frac{2}{3}\] multiply both sides by m \[f=\frac{2}{3}m\]

myininaya (myininaya):

now plug this into the second equation and solve for m

OpenStudy (anonymous):

5+1/3(2/3m) like that?

myininaya (myininaya):

\[\frac{2}{3}m=5+\frac{1}{3}(\frac{2}{3}m+m)\]

OpenStudy (anonymous):

thank you

myininaya (myininaya):

\[\frac{2}{3}m=5+\frac{1}{3}(\frac{2}{3}m+\frac{3}{3}m)\] \[\frac{2}{3}m=5+\frac{1}{3}(\frac{5}{3}m)\] \[\frac{2}{3}m=5+\frac{5}{9}m\] \[\frac{2}{3}m-\frac{5}{9}m=5\] \[\frac{6}{9}m-\frac{5}{9}m=5\] \[\frac{1}{9}m=5\] so m=45 right? but we want to know f+m so we need to know f too right?

myininaya (myininaya):

remember \[f=\frac{2}{3}m=\frac{2}{3}(45)=2(15)=30\]

myininaya (myininaya):

so f+m= what?